Supplement Question #1 Given the confidence interval (24.3, 29.5), find the margin of error and the sample mean. (5 pts)
Answer #1

The Confidence Interval is given by T ME

The Lower Limit = T - ME = 24.3 ----(1)

The Upper Limit = T + ME = 29.5 --- (2)

(a) Adding Equations (1) + (2), We get

2 * \bar{x} = 53.8

\bar{x} = 26.9

(b) Subtracting Equation (2) - (1)

2 * ME = 29.5 - 24.3 = 5.2

ME = 2.6